### v3

# Probabilistic Predictions of Process-Induced Deformation in Carbon/Epoxy Composites Using a Deep Operator Network

Abstract
Fiber reinforcement and polymer matrix respond differently to manufacturing conditions due to mismatch in coefficient of thermal expansion and matrix shrinkage during curing of thermosets. These heterogeneities generate residual stresses over multiple length scales, whose partial release leads to process-induced deformation (PID), requiring accurate prediction and mitigation via optimized non-isothermal cure cycles. This study considers a unidirectional AS4 carbon fiber/amine bi-functional epoxy prepreg and models PID using a two-mechanism framework that accounts for thermal expansion/shrinkage and cure shrinkage. The model is validated against manufacturing trials to identify initial and boundary conditions, then used to generate PID responses for a diverse set of non-isothermal cure cycles (time-temperature profiles). Building on this physics-based foundation, we develop a data-driven surrogate based on Deep Operator Networks (DeepONets). A DeepONet is trained on a dataset combining high-fidelity simulations with targeted experimental measurements of PID. We extend this to a Feature-wise Linear Modulation (FiLM) DeepONet, where branch-network features are modulated by external parameters, including the initial degree of cure, enabling prediction of time histories of degree of cure, viscosity, and deformation. Because experimental data are available only at limited time instances (for example, final deformation), we use transfer learning: simulation-trained trunk and branch networks are fixed and only the final layer is updated using measured final deformation. Finally, we augment the framework with Ensemble Kalman Inversion (EKI) to quantify uncertainty under experimental conditions and to support optimization of cure schedules for reduced PID in composites.
keywords:
Machine learning; Ensemble Kalman inversion; Deep operator networks; Residual stresses; Cure processes; Composite materials; Thermochemical modeling
1 Introduction
Thermoset fiber reinforced composites are widely adopted for lightweight structural applications in aerospace[57], wind energy [41], automotive[7, 8, 50], prosthetics, wearable electronics and sports equipment. A common route for their manufacturing is the lay-up of pre-impregnated sheets of unidirectional fibers and subsequently curing them under elevated pressure and temperature. During this process, residual stresses are induced due to the mismatch between the thermal expansion coefficients of the fiber and the matrix, as well as the shrinkage of the thermoset matrix as it cures and cross-links, thus affecting the mechanical properties of the manufactured component [62, 28, 66, 2, 14]. Release of some of these residual stresses due to material compliance and demolding results in process induced deformation (PID) that can render the manufactured part out of design tolerances and defective. Additional stresses may also be induced if the deformed and out-of-tolerance part is assembled and fastened in place, potentially leading to premature or catastrophic failure. Accounting for PID and designing the cure cycle for PID reduction is therefore a critical aspect of the design and manufacturing of composite parts. Residual stresses have been shown to reduce by as much as 25 % - 30 % through optimization of the cure cycle [63]. Several studies have tried to develop predictive capabilities using numerical methods. Constitutive laws have been proposed and demonstrated as part of finite element schemes to model cure dependent viscosity development and evolution of modulus and corresponding residual stresses [29, 9, 36, 67]. White et al. [62] investigated, both experimentally and analytically (on LamCure), the effect of changing the manufacturer recommended cure cycle on the resulting curvature for asymmetric cross-ply laminates. Li et al. [30] utilized a FE model for the multi-physics simulation to compute residual stresses, which was then modulated by an improved genetic algorithm to inversely determine the optimal process parameters to minimize residual stresses. Wang et al. [60] performed a fully coupled thermomechanical analysis to model the influence of cure cycles on residual stresses and degree of cure (DOC). They inferred that the temperature and duration of the second dwell had the greatest effect on residual stresses. Hui et al. [17] developed a collaborative multi-objective optimization strategy combining finite element-based cure process analysis with NGSA-II to demonstrate that the temperature gradient, residual stress, and process time can be reduced simultaneously.
Along with advances in physics-based modeling, rapid progress in machine learning (ML) has transformed manufacturing research by enabling efficient data-driven modeling, prediction, and optimization [3, 1, 44, 21]. Deep learning, in particular, has become a central tool across science and engineering. A growing body of work focuses on operator learning, which aims to approximate mappings between input and output functions, a setting that naturally arises in many scientific and engineering applications. For PID in composite structures, a range of ML architectures has been explored. Lin et al. [33] used neural networks, support vector regression, and $k$-nearest neighbors models to rapidly estimate differences in DoC from process parameters, while Fan et al. [11] employed convolutional neural networks to predict PID fields for laminates with varying stacking sequences. Theory-guided and simulation-informed approaches have also been proposed, including theory-guided ML with Gaussian process regression for process optimization [53], random forests for optimal cure design in co-cured composite structures [25], and simulation frameworks that capture competing mechanisms of residual stress evolution [32]. More recently, Liu et al. [35] combined viscoelastic finite element analysis, feature selection (Random Forest, LASSO), and artificial neural networks for rapid PID prediction, while Zhang et al. [65] introduced deep learning models tailored to complex thermo–chemo–mechanical interactions.
Operator learning provides a general ML framework for approximating mappings between function spaces rather than finite-dimensional vectors [38, 39, 54, 23, 34, 59, 4]. A prototypical architecture for this task is the Deep Operator Network (DeepONet) [38], which employs a branch–trunk structure, the branch network encodes the input function, and the trunk network parameterizes the dependence on spatial or temporal query points. Several alternative neural operator architectures have been proposed, including Fourier Neural Operators (FNO) [31] and U-Net-based operator models [52], which mainly differ in how they represent and propagate functional information. Physics-informed variants of DeepONet incorporate PDE residuals and other physical constraints into the loss function, enabling physically consistent predictions in data-scarce regimes [61, 20]. The Two-step DeepONet [26, 20, 49] extends this idea by composing multiple operator stages, thus increasing expressivity for complex multiscale or highly nonlinear mappings while retaining an operator-centric viewpoint.
Building on these developments, Feature-wise Linear Modulation (FiLM) [48] is extended to operator learning. FiLM was originally introduced for visual reasoning and has since been incorporated into architectures such as graph neural networks [5]. In the present work, a FiLM-augmented DeepONet surrogate model is employed to reconstruct the full histories of DoC, viscosity, and deformation, despite the fact that only the final deformation at the end of the cure cycle is available experimentally. To address the mismatch between the time discretization used in simulations and that available in experiments, a transfer learning strategy [43] is adopted, in which the DeepONet is first trained on high-fidelity simulation data and subsequently fine-tuned using experimental measurements through a loss function that enforces agreement between the predicted and measured final deformation.
While the FiLM-DeepONet framework provides accurate predictions for the evolution of DoC, viscosity, and deformation, performing reliable uncertainty quantification (UQ) for operator learning remains substantially more challenging. The associated inference problem is typically a very high-dimensional Bayesian inverse problem, which renders gradient-based sampling methods prohibitively expensive. Classical Bayesian approaches such as Hamiltonian Monte Carlo (HMC) [42] and the Laplace approximation [40] can, in principle, deliver high-quality uncertainty estimates, but they scale poorly, even when diagonal approximations of the Hessian are employed. As a practical alternative, Lakshminarayanan et al. [24] proposed deep ensembles trained with proper scoring rules, which can yield well-calibrated predictive distributions with only minor modifications to standard training pipelines. However, these ensemble methods have been criticized for lacking a clear Bayesian interpretation [64, 51, 46]. To help bridge this gap, Pearce et al. [45] introduced an ensemble scheme based on approximate Bayesian inference, in which model parameters are regularized around samples from a prior-like distribution, leading to more interpretable uncertainty estimates.
Another promising direction is ensemble Kalman inversion (EKI), a gradient-free ensemble method that has been successfully combined with UQ in scientific machine learning [13, 47]. Owing to its favorable scaling and rapid convergence, EKI is particularly well suited to high-dimensional operator-learning problems. In this work, we therefore employ EKI to train DeepONet while simultaneously obtaining uncertainty estimates. To further reduce the effective parameter dimension and improve computational efficiency, conventional multilayer perceptrons (MLPs) are replaced with Chebyshev enhanced Kolmogorov Arnold Networks (cKANs) within the DeepONet architecture. The cKANs employed here follow a modified DeepOKAN design, in which Chebyshev polynomials are incorporated into Kolmogorov Arnold Networks (KANs) [37, 55, 12] within a standard DeepONet framework. Inspired by Kolmogorov networks [56, 22], KANs are designed to adapt their activation patterns to the input data, thereby enabling more efficient forward evaluations and, consequently, faster generation of ensemble samples. The model developed as part of this study, thus couples operator learning with transfer learning by training a DeepONet surrogate on simulation data and refining it with experimental measurements to recover full deformation fields from limited observations, while EKI-based training simultaneously yields uncertainty estimates, providing experimental predictions that are both accurate and uncertainty-aware.
This paper is organized as follows. In Section 2, the experimental materials, methods, and manufacturing trials are presented. Building on this foundation, Section 3 details the simulation framework and datasets used to model the cure process. Section 4 then introduces the operator-learning perspective, presenting the DeepONet and FiLM-DeepONet architectures, together with the associated network design and training strategy. The resulting predictive capabilities for cure processing are assessed in Section 5. To quantify uncertainty in these predictions, Section 6 develops DeepONet ensembles for epistemic uncertainty and incorporates the EKI methodology within the DeepONet-based framework. Leveraging these ingredients, Section 7 focuses on the optimization of the cure temperature profile. Finally, Section 8 summarizes the main findings and outlines directions for future research.
2 Experimental Materials and Methods
This section describes the experimental setup, specimen preparation, and materials used in this study.
2.1 Experimental Approach
Asymmetric laminates inherently develop residual stresses during cure, resulting in measurable out-of-plane deformation at the end of the process [62]. To leverage this effect, an asymmetric $[0/90]$ cross-ply laminate configuration was selected. This layup provides a clear and reproducible manifestation of PID, enabling direct comparison between simulations and experiments. The simulation framework used in this study captures the physical mechanisms governing residual stress evolution and the resulting PID. Using this framework, two optimized non-isothermal cure cycles were identified. Experiments were then conducted in which the baseline cure cycle, along with the two optimized cycles, were implemented on a laboratory-scale setup, and the resulting PID was measured. These experiments serve two purposes: (i) to validate the simulation model by confirming that the initial and boundary conditions correctly represent the physical process, and (ii) to provide experimental data that complement the larger simulation-generated dataset for DeepONet model development.
The laminate layup was carried out with particular care to ensure consistent ply orientation across all specimens, thereby reducing variability associated with fiber misalignment. Samples fabricated from the selected thermoset prepreg were cured under constant consolidation pressure using one baseline isothermal cycle and two optimized non-isothermal cycles aimed at mitigating PID. The temperature histories of the specimens were recorded in real time and subsequently used to predict the deformation history using a FiLM-enhanced DeepONet.
2.2 Materials and Methods
[…]
$\alpha=\left(\frac{\Delta H_{\text{residual cure}}}{\Delta H_{\text{full cure}}}\right)\times 100,$ (1)
where:
  • 1.
    $\Delta H_{\text{residual cure}}$ = Enthalpy of reaction of the as-received B-stage resin,
  • 2.
    $\Delta H_{\text{full cure}}$ = Ultimate enthalpy of reaction of freshly mixed resin prior to prepreg manufacturing and use.
Two differential scanning calorimetry (DSC) methods are typically used to determine $\alpha_{i}$: (1) measuring the residual heat of reaction in the as-received material, or (2) assessing the shift in the glass transition temperature ($T_{g}$) [58]. In this study, the first method was adopted, wherein the residual heat of reaction was measured to determine $\alpha_{i}$, following the approach described by Sun et al. [58]. Hargis et al. [15] performed a detailed calorimetric analysis of AS4/3501-6 using DSC to characterize the cure kinetics. Since cure kinetics generally describe the behavior of the thermoset epoxy resin matrix, it is important to accurately determine the resin mass fraction in the prepreg when DSC samples are taken directly from fiber-reinforced prepregs [19]. A more common alternative is to obtain the thermoset epoxy system from the manufacturer as separate, unmixed components, mix them immediately prior to the DSC run, and perform DSC on the neat epoxy system [27, 6]. However, this latter approach yields the total heat of reaction for freshly mixed resin and may not accurately represent the state of the epoxy in the prepreg, where some prereaction has already taken place. Consequently, several studies have instead performed DSC measurements directly on AS4/3501-6 prepreg and calculated the heat of reaction while compensating for the fiber mass fraction present in the DSC sample [19, 6]. The values reported in the literature are summarized in Table 1.
Reference Year Material Composition Measurement Conditions Enthalpy of Reaction $H_{t}$ (J/g) Details on Pre-reaction
Lee et al. [27] 1982 3501-6 Resin @ 20 °C/min to 326.85 °C $473.6\pm 5.4$ NA
Hou et al. [16] 1988 3501-6 Resin @ 20 °C/min $502\pm 21$ NA
White et al. [63] 1993 AS4 CF/3501-6 Prepreg (42% resin mass) @ 20 °C/min to 350 °C 182.87 (prepreg) 435 (comp. for resin) NA
Kim et al. [19] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 20 °C/min to 350 °C 433.7 (comp. for resin) NA
Kim et al. [19] 2002 AS4 CF/3501-6 Prepreg (37.4% resin mass) @ 2 °C/min to 350 °C 456.7 (comp. for resin) NA
Chern et al. [6] 2002 3501-6 Resin @ 20 °C/min $508\pm 19$ 2% cure (manufacturer’s note)
Hargis et al. [15] 2006 3501-6 Resin @ 20 °C/min $382.5\pm 20$ Significant pre-reaction reported
Table 1: Comparison of the enthalpy of reaction ($H_{t}$) for the 3501-6 resin system and AS4 CF/3501-6 epoxy prepreg under dynamic DSC conditions. The measurement method (Dynamic) is common to all studies.
Based on the values reported in Table 1 and the procedure in Equation (1), $\mathrm{DoC}_{0}$ was estimated to lie within the range of 17.68%–31.21% when the measurement standard deviations were taken into account. Using the reported mean values, $\mathrm{DoC}_{0}$ was calculated to be 24.7%. Accordingly, an initial condition of $\mathrm{DoC}_{0}=0.3$ (30%) was assumed for the simulations. This approximation also accounts for additional curing that may have occurred during the several months between prepreg manufacture and its use in the experiments.
The material used in the compression molding trials was characterized under comparable DSC conditions using a Netzsch Polyma 214 DSC (Netzsch Instruments Inc., Burlington, MA, USA), with nitrogen purge gas and an empty aluminum pan as reference. The results from the first dynamic heating run at 20 °C/min are presented in Figure 1 and show good agreement with the prepreg values reported by White et al. [63] (see Table 1).
Figure 1: Dynamic DSC run at heating rate of 20 °C/min indicating heat flow versus time for the AS4 Cf/3501-6 epoxy prepreg, with the area under the peaks indicating the heat of reaction resulting from the curing of the resin due to heat input during the DSC run
2.3 Manufacturing Trials
The experimental validation runs were performed on a lab-scale setup for compression molding. The setup utilized a 10 kN load capacity Instron 68TM-10 (Instron,Norwood, MA, USA) servo-electric universal testing machine (UTM) with a 10 kN load cell and a temperature chamber 3119-615 (Buckinghamshire, England) with a temperature range of -50 °C to 450 °C, capable of executing programmed heating and cooling cycles. The setup is as illustrated in Figure 2.
Figure 2: Experimental setup to consolidate and heat unbalanced [0/90] lay-up of prepregs to replicate the cure cycle to be followed during manufacturing, with a universal testing machine for precise displacement control of the compression plates, a 10 kN load cell measure and control the consolidfation force applied, and K-type thermocouples in contact with the material to measure and log temperature change.
Figure 3: Specimen preparation methodology indicating layup of a 150 mm (6”) x 150 mm (6”) unbalance laminate followed by cutting into 6 specimens of 25.4 mm (1”) width, arrangement of the specimens on the compression plates in contact with the thermocouples to enable in-situ temperature measurement and logging
The unbalanced ply layup of 152 mm (6”) x 152 mm (6”) unidirectional prepreg squares, were cut into 152 mm (6”) x 25.4 mm (1”) rectangular specimens, and placed onto a 3.17 mm (1/8th inch) thick aluminum caul plate. K-type thermocouples, with a measurement range of -50 °C to 400 °C, were placed in contact with the samples to accurately measure the material temperature. Additional thermocouples were also placed in contact with the top and bottom compression plates and the thermal chamber to ensure the thermal cycle was executed as programmed. Initial trials helped determine the correct heating rates, cooling rates and temperature setpoints to be programmed, so that the material was subjected to the right heating and cooling rates. Thus, the difference associated with the dynamic temperature lag between the chamber temperature and material temperature was accounted for. The specimen preparation is illustrated in Figure 3. Henkel Loctite Frekote 55-NC, a semi-permanent, non-contaminating mold release agent was applied onto the compression plates to enable easy release of the cure laminates after curing.
2.4 Experimental Cure Cycles and Measurements
[…]
Figure 4: Temperature–time profiles for the three cure cycles: an isothermal baseline cycle and two non-isothermal cycles identified through the optimization scheme in [32]. The curves show the measured specimen temperature, the thermal chamber environment, and the corresponding simulated cure profiles.
Figure 5: Measurement of process induced deformation in unbalanced ply lay-up. The chord center point was identified and the perpendicular distance from that point to the curved unbalance laminate was measured
Thermal cycle Run Specimen 1 PID (mm) Specimen 2 PID (mm) Specimen 3 PID (mm) Specimen 4 PID (mm) Mean PID (mm) Standard deviation (mm) Simulated PID (mm)
Baseline 1 41.5 35.0 37.0 41.5 38.75 3.278 39.23
Baseline 2 39.0 36.5 34.5 36.667 2.254 39.23
Baseline 3 36.0 41.0 36.0 36.5 37.375 2.428 39.23
Optimal R11 1 36.5 37.0 37.5 37.000 0.500 36.64
Optimal R11 2 33.5 35.0 35.0 34.500 0.866 36.64
Optimal R11 3 31.0 33.5 36.0 33.500 2.500 36.64
Optimal R21 1 31.0 32.0 35.5 32.833 2.362 36.84
Table 2: Comparison of experimentally measured and simulated PID (mm) for the three cure cycles: an isothermal baseline cycle and two non-isothermal cycles identified through the optimization scheme in [32].
3 Model and Simulations
A comprehensive study by Limaye et al. [32] demonstrated that the cure cycle governs phase transitions in polymer laminates, progressing from viscous to viscoelastic and ultimately to elastic behavior. The study further highlighted the critical role of thermo-mechanical interactions in modulus evolution and the development of residual stresses and deformations. These coupled phenomena were represented using Polymerization–Gelation–Vitrification (PGV) plots, which characterize material behavior as a function of the applied cure cycle. To capture these complex interactions, a computational cure analysis framework was developed using ABAQUS and COMPRO in a sequential two-step approach, comprising thermo-chemical analysis followed by stress–deformation analysis. The model predicted temperature evolution, DoC, and laminate deformation. These predictions were validated against existing literature, showing strong agreement in both DoC profiles and deformation measurements.
An optimization scheme was developed and implemented for the R11 and R21 cases using the NSGA-II genetic algorithm, with the objective of minimizing PID. This was accomplished by adjusting the slopes of the cure cycle’s two linear ramp segments, resulting in non-isothermal cure cycles that reduced overall deformation by balancing cure shrinkage with thermal expansion during processing. For these cases, $\mathrm{DoC}_{0}$ was prescribed, while the heating time was fixed at 1 minute for R11 and 10 minutes for R21. The optimization was performed subject to the constraints of achieving a high final cure state ($\mathrm{DoC}\geq 0.990$) and promoting interaction between thermal expansion and chemical shrinkage effects.
In the present work, the cure model is regenerated to evaluate non-isothermal cure cycles while accounting for material pre-cure through the prescribed value of $\mathrm{DoC}_{0}$. The end-of-cycle deformation predicted for the simulated cure cycles R11 and R21 is compared against experimental measurements (see Table 2 and Figure 4) to validate the model. The corresponding simulation data are subsequently used to assess the performance of the transfer learning approach. Because the final-time deformation is available experimentally, the R11 and R21 cases serve as validation benchmarks for the FiLM-DeepONet model to evaluate the accuracy of deformation-history predictions enabled by transfer learning.
As shown in Figure 6(a), a family of simulated temperature profiles is generated by varying an intermediate temperature coordinate $A=(t_{1},T_{1})$ between the fixed start point $(t_{0},T_{\mathrm{start}})=(0.333,20)$ and peak point $(t_{2},T_{\mathrm{peak}})=(171.658,179.905)$, while keeping the end point $(t_{3},T_{\mathrm{end}})=(205,20)$ fixed. This intermediate coordinate controls both the duration and the temperature level of the dwell during the heating stage. Figure 6(b) shows the corresponding histories of DoC, viscosity, and deformation for different choices of the temperature coordinate $A$, illustrating the influence of $\mathrm{DoC}_{0}$ on the deformation evolution.
(a) Parametrization of the cure temperature profile with an intermediate control point $A=(t_{1},T_{1})$ and fixed start $(t_{0}=0.333,\;T_{\text{start}}=20.000)$, peak $(t_{2}=171.658,\;T_{\text{peak}}=179.905)$, and end $(t_{3}=205.000,\;T_{\text{end}}=20.000)$ temperatures, together with imposed temperature histories at selected temperature coordinates $T_{c}$ for cure cycles used to generate the simulation data in ABAQUS.
(b) DoC and viscosity profiles from the thermo–chemical simulation, together with the corresponding deformation from the stress–deformation simulation, for $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$.
Figure 6: Simulation data for a laminate subjected to a cure cycle. The top panel combines a parametrized cure temperature profile with an intermediate control point $A=(t_{1},T_{1})$ and fixed start, peak, and end temperatures, and the corresponding imposed temperature histories at different temperature coordinates $T_{c}$ used in simulation data. The bottom panel shows the simulated evolution of DoC, viscosity, and deformation for two initial degrees of cure, $\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$.
4 Operator Learning
Operator learning provides a machine learning framework for approximating mappings between function spaces rather than finite-dimensional vectors [38, 39, 54, 23, 34]. The goal is to learn solution operators that map input functions, such as initial or boundary conditions, to output functions, for example the corresponding solutions of partial differential equations. A prototypical architecture for this task is DeepONet [38], which employs a branch–trunk structure, the branch network encodes the input function, while the trunk network parameterizes the dependence on spatial or temporal query points.
4.1 DeepONet
DeepONet approximates nonlinear operators between function spaces using two neural networks in parallel, a branch network and a trunk network. In its general formulation, the branch network encodes the input function $f$ by sampling it at a fixed set of sensor locations $\{\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{k}\}$. The sampled values,
$\big(f(\bm{x}_{1}),f(\bm{x}_{2}),\ldots,f(\bm{x}_{k})\big),$
are mapped through the branch network to produce a latent feature vector $\mathbf{b}\in\mathbb{R}^{p}$. For a batch of $m$ input functions, this results in a matrix $B\in\mathbb{R}^{m\times p}$, with each row corresponding to one input sample. The trunk network receives as input a spatial or temporal query point $\bm{x}\in\mathbb{R}^{d}$ at which the operator output is to be evaluated and produces a corresponding latent representation $\bm{\phi}(\bm{x})\in\mathbb{R}^{p}$. For $n$ evaluation points, the trunk network outputs a matrix $\Phi\in\mathbb{R}^{n\times p}$. The operator output $\mathcal{G}(f_{i})$ evaluated at a query point $\bm{x}_{j}$ is obtained through the dot–product contraction of the branch and trunk outputs:
$\mathcal{G}(f_{i})(\bm{x}_{j})\approx\sum_{l=1}^{p}B_{il}(f_{i};\theta_{B})\,\Phi_{jl}(\bm{x}_{j};\theta_{T}),$ (2)
where $\theta_{B}$ and $\theta_{T}$ denote the trainable parameters of the branch and trunk networks, respectively.
In this work, DeepONet is configured to learn the mapping from the imposed temperature profile $\mathbf{T}$ of a cure cycle to the DoC, viscosity, and deformation history of the laminate. The branch network takes the temperature field $\mathbf{T}$ as input, represented by samples of the cure-cycle temperature profile for different temperature coordinates $T_{c}$ (see Figure 6(a)),
$T_{c}=\{T(t_{1}),T(t_{2}),\ldots,T(t_{k})\},$
which serve as a discretized representation of the input function. The trunk network takes the query time $t$ as input and, for a set of evaluation times $\{t_{1},t_{2},\ldots,t_{n}\}$, produces the corresponding latent basis functions $\Phi(t_{j})$, from which the time-dependent outputs are constructed.
4.2 FiLM-enhanced DeepONet
As described in Section 4.1, the standard DeepONet architecture learns an operator by combining a branch network, which encodes the input function, with a trunk network, which represents the evaluation coordinates. However, in the present problem, the laminate response depends not only on the applied temperature history but also on $\mathrm{DoC}_{0}$. As illustrated in Figure 6(b), variations in $\mathrm{DoC}_{0}$ significantly influence the evolution of $\mathrm{DoC}$, viscosity, and deformation.
To incorporate this additional physical dependence, we augment the branch network with Feature-wise Linear Modulation (FiLM), resulting in a DeepONet architecture. FiLM provides a mechanism for conditioning intermediate branch features on an auxiliary variable—in this case, $DoC_{0}$—allowing the operator representation to adapt dynamically to different initial material states. For a hidden-layer activation $h$ in the branch network, the FiLM transformation is given by
$h^{\prime}=\gamma(DoC_{0})\,h+\beta(DoC_{0}),$
where $\gamma(DoC_{0})$ and $\beta(DoC_{0})$ are learned affine functions of the conditioning input. More generally, for a vector of hidden features $\mathbf{h}$ and conditioning input $\mathbf{c}$,
$\text{FiLM}(\mathbf{h}\mid\mathbf{c})=\gamma(\mathbf{c})\odot\mathbf{h}+\beta(\mathbf{c}),$
with $\odot$ denoting element-wise multiplication. Through this modulation, the branch network produces a representation of the temperature history that is explicitly conditioned on $\mathrm{DoC}_{0}$.
In the Feature-wise Linear Modulation DeepONet used here, the branch network receives the sampled temperature profile $\mathbf{T}$ together with $DoC_{0}$. Its output is partitioned into three latent vectors corresponding to the three predicted quantities:
$\mathbf{h}_{\text{branch}}(\mathbf{T},d_{0};\theta_{\text{branch}})=\big[\,\mathbf{h}_{d},\;\mathbf{h}_{v},\;\mathbf{h}_{\varepsilon}\,\big]\in\mathbb{R}^{3G},$
where $\mathbf{h}_{d}$, $\mathbf{h}_{v}$, and $\mathbf{h}_{\varepsilon}\in\mathbb{R}^{G}$ encode the operator representations for DoC, viscosity, and deformation, respectively. The trunk network takes the query time $t$ as input and produces a shared latent basis:
$\mathbf{h}_{\text{trunk}}(t;\theta_{\text{trunk}})\in\mathbb{R}^{G}.$
Following the DeepONet formulation, the predicted outputs are obtained by inner products between the FiLM-modulated branch features and the trunk basis:
$\hat{d}(t)=\mathbf{h}_{d}^{\top}\mathbf{h}_{\text{trunk}}(t),\qquad\hat{v}(t)=\mathbf{h}_{v}^{\top}\mathbf{h}_{\text{trunk}}(t),\qquad\hat{\varepsilon}(t)=\mathbf{h}_{\varepsilon}^{\top}\mathbf{h}_{\text{trunk}}(t).$
In this study, the DeepONet architecture is enhanced using Feature-wise Linear Modulation (FiLM). For simplicity, we refer to this FiLM-augmented model as DeepONet throughout the remainder of the paper.
5 DeepONet-Based Framework for Modeling Cure-Induced Material Response
Figure 7 presents a schematic of the DeepONet architecture used to predict the DoC, viscosity, and deformation of the composite during the cure process. In the first stage, DeepONet is trained on simulation data, the branch network receives the temperature profile, the trunk network takes time as input, and their combination yields the full temporal histories of DoC, viscosity, and deformation. In the second stage, the trained DeepONet is employed as a surrogate model for the experiments. The simulated temperature profiles are defined on a common time interval from 0 to $205$ minutes, whereas the experimental temperature profiles span different time windows (e.g., up to approximately $255$, $211.6$, or $227.6$ minutes, depending on the test) and are available only in limited quantity. Training a new network solely on this experimental dataset would be inefficient and prone to overfitting. Instead, a transfer-learning strategy is adopted, the weights learned from the simulation stage are retained, and only the last layer of the branch network is updated so that the predicted deformation matches the experimentally observed final deformation. An additional loss term penalizes the discrepancy between the predicted and measured deformation at the final time, yielding corrected deformation histories for each experimental temperature profile while still leveraging the rich operator learned from the simulations.
Figure 7: Schematic of the DeepONet architecture for predicting DoC, viscosity, and deformation of a composite material during the cure process. In the first stage, DeepONet is trained on simulation data, the branch network takes the temperature history as input, while the trunk network takes time as input to predict the DoC, viscosity, and deformation histories. In the second stage, the trained DeepONet is used as a surrogate model, and transfer learning is applied by modifying the last layer of the branch network to satisfy the final deformation observed in the experimental data. An additional loss term is introduced to compare the predicted deformation with the experimental deformation, resulting in a final prediction of the deformation history under the experimental temperature profile.
As illustrated in Figure 7, the branch network takes as input the discretized temperature field $\mathbf{T}$, sampled at a set of temperature coordinates, together with $\mathrm{DoC}_{0}$, and outputs three sets of latent coefficients corresponding to the histories of DoC, viscosity, and deformation. The trunk network receives the time coordinates and produces a corresponding set of basis functions; the predicted histories are then obtained by combining the branch and trunk outputs. Both branch and trunk networks employ three hidden layers with 20 neurons each and hyperbolic tangent ($\tanh$) activation functions, followed by an output layer with 20 neurons. The model is trained on simulation data using the Adam optimizer with an exponentially decaying learning rate, initialized at $10^{-3}$ and updated according to an ExponentialDecayschedule, for up to $10^{5}$ iterations with early stopping.
Figure 8: Comparison of FiLM-DeepONet predictions with true simulation results at three temperature coordinates, including DoC for two initial conditions ($\mathrm{DoC}_{0}=0.3$ and $\mathrm{DoC}_{0}=0.001$), viscosity, and deformation. $\mathrm{DoC}_{0}$ is provided as a conditioning input to the branch network.
Figure 8 compares DeepONet predictions with reference simulation results for $\mathrm{DoC}$, viscosity, and deformation over time under four temperature profiles, using two initial conditions ($\mathrm{DoC}_{0}=0.001$ and $\mathrm{DoC}_{0}=0.3$). In the FiLM-enhanced architecture, $\mathrm{DoC}_{0}$ conditions the branch network, enabling the model to capture how the coupled thermochemical and mechanical response varies with the initial cure state. The close agreement between predictions and reference data demonstrates both the accuracy and generalization capability of the trained DeepONet. The rightmost panel further indicates that $\mathrm{DoC}_{0}$ strongly influences the final deformation pattern, justifying its inclusion as a conditioning variable. As shown in the figure, viscosity grows approximately exponentially with time, leading to steep gradients and a wide dynamic range. Such highly skewed data can be challenging for neural networks to learn directly. To mitigate this, viscosity was log-transformed prior to training,
$\mu_{\text{log}}=\log(\mu+\varepsilon),$
where $\varepsilon=10^{-8}$ is a small constant introduced for numerical stability. The figure also illustrates that FiLM enables the network to clearly distinguish responses corresponding to different initial degrees of cure.
6 Uncertainty Quantification for DeepONet
UQ is crucial for assessing the reliability of DeepONet predictions, particularly with noisy or limited data. By explicitly characterizing confidence in the model outputs, UQ enables more robust and trustworthy deployment of operator-learning frameworks in practical applications [18]. In this section, we present UQ results for DeepONet, focusing on epistemic uncertainty and EKI. We demonstrate how transfer learning under uncertainty can predict deformation histories, together with associated uncertainty bounds, for the experimental temperature profiles.
6.1 DeepONet Ensembles for Epistemic Uncertainty
To quantify epistemic uncertainty, we train an ensemble of DeepONet models using different random seeds for weight initialization, while keeping the training/validation split fixed. For each input, this ensemble yields multiple predictions, from which we compute the mean and standard deviation across seeds as a measure of epistemic uncertainty. In practice, we observe that the ensemble spread is relatively small, as independently initialized networks trained on the same data tend to converge to similar solutions. After this initial training, the ensemble is further refined via transfer learning by augmenting the loss function with a term that penalizes the mismatch between the predicted deformation at the final time step and the final deformation value reported in the experimental data. This encourages the network to remain consistent with the full deformation profile while accurately matching the terminal deformation. Figure 9 (left) illustrates the deformation histories for two baseline specimen experiments. The curves show the ensemble-averaged deformation profile (mean over multiple seeds), with an uncertainty band corresponding to the ensemble standard deviation. The experimental final deformation is indicated as a single marker, demonstrating that the transfer-learned predictions closely match the reported terminal values. Figure 9 (right) shows the predicted deformation history for the optimal R21 case obtained from the transfer-learned DeepONet ensemble, together with the corresponding simulation results used for validation. Since the full deformation history is not available from the experiments, the ensemble predictions are evaluated against the simulated deformation profile. The shaded region denotes the standard deviation across different seeds, providing a measure of the epistemic uncertainty associated with the DeepONet ensemble.
Figure 9: Mean and standard deviation of DeepONet ensemble predictions for deformation. (Left) Deformation histories for two baseline specimen experiments. The solid curves show the ensemble-averaged predictions (mean over six independently trained models with different random seeds), and the shaded bands indicate the corresponding standard deviation. The experimental final deformation for each experiment is shown as a single marker, illustrating the agreement between the transfer-learned predictions and the measured terminal values. (Right) Deformation history for optimal R21, comparing the transfer-learned ensemble prediction against the corresponding simulation data. Here, the shaded region again denotes the standard deviation across seeds, visualizing the epistemic uncertainty of the DeepONet ensemble.
6.2 Ensemble Kalman Inversion
The ensemble Kalman filter (EnKF), originally introduced by Evensen in 1994 for data assimilation in time dependent dynamical systems, has become a widely used tool owing to its robustness, ease of implementation, and strong numerical performance [10]. Building on this foundation, iterative ensemble Kalman methods have been developed for solving inverse problems. In this context, the inferred solution is constrained to the subspace spanned by the initial ensemble, which allows for error estimates relative to the best approximation achievable within that subspace. Numerical studies have shown that this ensemble-based, derivative free strategy can attain accuracy comparable to classical least squares approaches, and in many cases close to the optimal approximation.
EKI is a gradient free methodology specifically designed for inverse problems. Recent work [47] applied EKI to the training of DeepONets and demonstrated that it achieves substantially faster convergence while providing more reliable uncertainty quantification than sampling based alternatives such as HMC. Motivated by these results, we integrate EKI into our operator learning framework to obtain efficient and informative uncertainty estimates for DeepONet.
(a) Comparison of ensemble predictions for DoC, viscosity, and deformation obtained with EKI against the corresponding simulation data.
(b) Ensemble sample trajectories for DoC, viscosity, and deformation, with solid lines indicating the reference simulation responses.
Figure 10: Prediction uncertainty and ensemble behavior for DoC, viscosity, and deformation using EKI. (a) Uncertainty bands showing ensemble mean predictions (solid lines), reference simulation responses (dashed lines), and shaded regions corresponding to one standard deviation of the ensemble. (b) Ensemble trajectory plots illustrating the spread of individual samples around the reference simulation responses (solid lines).
In detail, we assume that the noisy observations are generated by
$G_{i}^{(l)}=\mathcal{G}^{\star}(u^{(l)})(y_{i}^{(l)})+\varepsilon_{i}^{(l)},\qquad\varepsilon_{i}^{(l)}\sim\mathcal{N}(0,\sigma_{l}^{2}I),$
where $\mathcal{G}^{\star}$ denotes the (unknown) ground-truth operator and $G_{i}^{(l)}$ is the noisy observation at the $i$-th query location for the $l$-th input function. The noise variables $\varepsilon_{i}^{(l)}$ are independent zero-mean Gaussian with noise level $\sigma_{l}$. For the $l$-th input function $u^{(l)}$ we collect the corresponding noisy outputs in $\mathcal{D}^{(l)}=\{G^{(l)}_{i}\}_{i=1}^{P}$, where $G^{(l)}_{i}$ denotes the noisy output at $P$ query locations. By stacking the datasets for all $N$ inputs we obtain the full noisy dataset $\mathcal{D}=\{\mathcal{D}^{(l)}\}_{l=1}^{N}$. We assume that all observations are independent and identically distributed (i.i.d.) and follow a Gaussian likelihood. To approximate the true operator we introduce a DeepONet operator $\mathcal{G}_{\bm{\theta}}\approx\mathcal{G}^{\star}$ parameterized by $\bm{\theta}\in\mathbb{R}^{N_{\theta}}$. The operator learning task then becomes a Bayesian inverse problem with likelihood
$p(\mathcal{D}\mid\bm{\theta})=\prod_{l=1}^{N}\prod_{i=1}^{P}\frac{1}{\sqrt{2\pi\sigma_{l}^{2}}}\exp\left(-\,\frac{\bigl(G^{(l)}_{i}-\mathcal{G}_{\bm{\theta}}(u^{(l)})(y^{(l)}_{i})\bigr)^{2}}{2\sigma_{l}^{2}}\right).$ (3)
Assuming a standard multivariate Gaussian prior on $\bm{\theta}$ and applying Bayes’ formula, the posterior distribution is given by
$p(\bm{\theta}\mid\mathcal{D})\propto p(\mathcal{D}\mid\bm{\theta})\,p(\bm{\theta}).$
To approximate this posterior distribution, EKI considers evolving an ensemble of particles according to the dynamical system
$\begin{split}\bm{\theta}_{t}&=\bm{\theta}_{t-1}+\bm{\xi}_{t},\qquad\bm{\xi}_{t}\sim\mathcal{N}(0,Q),\\[2.84526pt] \mathbf{y}_{t}&=\mathcal{F}(\bm{\theta}_{t})+\bm{\zeta}_{t},\qquad\bm{\zeta}_{t}\sim\mathcal{N}(0,R),\end{split}$
where $\bm{\xi}_{t}$ is an artificial process noise with covariance matrix $Q\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$, and $\mathcal{F}(\bm{\theta}_{t})$ denotes the forward map obtained by evaluating the DeepONet operator on the fixed training inputs,
$\mathcal{F}(\bm{\theta}_{t})=\mathrm{vec}\Bigl(\bigl\{\mathcal{G}_{\bm{\theta}_{t}}(u^{(l)})(y_{i}^{(l)})\bigr\}_{l=1,\dots,N;\;i=1,\dots,P}\Bigr)\in\mathbb{R}^{NP},$
while $\bm{\zeta}_{t}$ represents the (flattened) observation noise with covariance $R\in\mathbb{R}^{NP\times NP}$. Given an ensemble of initial particles $\{\bm{\theta}_{0}^{(j)}\}_{j=1}^{J}$ and using the marginals of the joint Gaussian distribution, the EKI update for the ensemble $\{\bm{\theta}_{t}^{(j)}\}_{j=1}^{J}$ reads
$\hat{\bm{\theta}}_{t}^{(j)}$ $=\bm{\theta}_{t-1}^{(j)}+\bm{\xi}_{t}^{(j)},$ $\bm{\xi}_{t}^{(j)}$ $\sim\mathcal{N}(0,Q),$ (2.10)
$\hat{\mathbf{y}}_{t}^{(j)}$ $=\mathcal{F}\bigl(\hat{\bm{\theta}}_{t}^{(j)}\bigr),$ (2.11)
$\bm{\theta}_{t}^{(j)}$ $=\hat{\bm{\theta}}_{t}^{(j)}+C_{t}^{\hat{\theta}y}\bigl(C_{t}^{yy}+R\bigr)^{-1}\bigl(\mathbf{y}-\hat{\mathbf{y}}_{t}^{(j)}+\bm{\zeta}_{t}^{(j)}\bigr),$ $\bm{\zeta}_{t}^{(j)}$ $\sim\mathcal{N}(0,R),$ (2.12)
where $C_{t}^{yy}$ and $C_{t}^{\hat{\theta}y}$ are the empirical covariance matrices, defined as
$C_{t}^{yy}$ $=Y_{t}Y_{t}^{T},C_{t}^{\hat{\theta}y}=\Theta_{t}Y_{t}^{T},$ (2.13)
$Y_{t}$ $=\frac{1}{\sqrt{J-1}}\bigl[\hat{\mathbf{y}}_{t}^{(1)}-\bar{\mathbf{y}}_{t},\,\ldots,\,\hat{\mathbf{y}}_{t}^{(J)}-\bar{\mathbf{y}}_{t}\bigr],$ (2.14)
$\Theta_{t}$ $=\frac{1}{\sqrt{J-1}}\bigl[\hat{\bm{\theta}}_{t}^{(1)}-\bar{\bm{\theta}}_{t},\,\ldots,\,\hat{\bm{\theta}}_{t}^{(J)}-\bar{\bm{\theta}}_{t}\bigr].$ (2.15)
Here, $\bar{\bm{\theta}}_{t}$ and $\bar{\mathbf{y}}_{t}$ denote the sample means of the prior ensembles $\{\hat{\bm{\theta}}_{t}^{(j)}\}_{j=1}^{J}$ and $\{\hat{\mathbf{y}}_{t}^{(j)}\}_{j=1}^{J}$ at iteration $t$, respectively.
Predictions without transfer learning Predictions with transfer learning
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Ensemble samples generated using EKI for DoC, viscosity, and deformation based on simulation data. Solid lines denote the mean predictions and shaded regions indicate the corresponding standard deviation. Results are shown for two experimental baseline specimens; the left panels present predictions without transfer learning, while the right panels show predictions with transfer learning. Blue dots indicate the final deformation measured from each experimental dataset.
In detail, we use an ensemble of size $J=2000$ and run EKI for 1000 iterations. Both the input and output data are perturbed with $1\%$ additive Gaussian noise, leading to an observation covariance of $R=0.01I$. To mitigate ensemble collapse, we also introduce artificial process noise with covariance $Q=0.002I$. The trunk and branch networks each contain two hidden layers with 10 neurons per layer, and their output layers produce 30-dimensional feature vectors. This compact network design substantially reduces the total number of trainable parameters.
Figure 10 summarizes the resulting predictions of DoC, viscosity, and deformation on the simulation test dataset together with their associated uncertainty estimates. Panel (a) displays uncertainty bands, where solid lines denote the ensemble mean, dashed lines indicate the reference simulation responses, and shaded regions correspond to one standard deviation of the ensemble. Panel (b) shows ensemble “spaghetti” trajectories, illustrating the spread of individual samples around the reference responses (solid lines). We observe that the prediction accuracy is high and that the confidence intervals consistently cover the ground truth trajectories, indicating that the proposed EKI–DeepONet framework yields meaningful and well-calibrated uncertainty quantification.
For inference on the experimental data, we employ a transfer learning strategy. Because the experimental loss function is defined only at the final time point, the risk of overfitting is particularly high. To mitigate this, we fine-tune only the parameters in the last layer of the branch network, keeping all other parameters fixed at the values obtained from training on the simulation data. Moreover, we use Tikhonov-regularized EKI with a regularization parameter of $0.1$ to further control overfitting. Figures 11 and 12 show the predicted deformation on the experimental dataset, before and after transfer learning, for the two case studies (two baseline specimens). After transfer learning, the prediction at the final time is significantly closer to the experimental ground truth, while the uncertainty along the earlier portions of the trajectory increases, resulting in more conservative and better-posed predictive uncertainty.
Figure 12: Ensemble samples generated using EKI for the deformation in experimental optimal cases R11 and R21, shown alongside the corresponding simulation results for comparison. The solid black line represents the simulation data used for validation, while the colored lines denote the ensemble samples generated by EKI.
Figure 12 shows ensemble samples generated using EKI for the deformation in experimental cases R11 and R21, plotted together with the corresponding simulation results for comparison. This figure illustrates the ability of the network to generate predictive sample trajectories. To assess performance, we validate the ensemble predictions against the simulation data for cases R11 and R21. The close agreement between the simulation curves and the EKI-generated ensembles indicates that the method provides a reasonable match for these cases.
7 Optimization of the cure temperature profile
As illustrated in Figure 6, the cure process is parameterized by the time and temperature of an intermediate point $A=(t_{1},T_{1})$, which determines how long and at what temperature the resin is heated during the dwell stage. In the optimization problem, the start, peak, and end temperatures and their corresponding times are fixed. The sole design variable is the intermediate point $A=(t_{1},T_{1})$. By appropriately choosing $(t_{1},T_{1})$, the intermediate heating segment of the cure cycle is reshaped to reduce the final process induced deformation while still achieving full cure.
Let $DoC(t;T)$ denote the DoC and $u(t;T)$ the magnitude of process-induced deformation predicted by the trained DeepONet surrogate under a given temperature schedule $T(\cdot)$. The optimization problem is
$\min_{t_{1},T_{1}}$ $J(t_{1},T_{1})=\big\|u\!\left(t_{3};\,T(\cdot;t_{1},T_{1})\right)\big\|,$ (4)
s.t. $\text{DoC}\!\left(t_{3};\,T(\cdot;t_{1},T_{1})\right)\;\geq 0990\quad\text{(full cure)},$
$t_{0}+\Delta t\;\leq\;t_{1}\;\leq\;t_{2}-\Delta t,$
$T_{\text{start}}\;\leq\;T_{1}\;\leq\;T_{\text{peak}},$
$m_{1}(t_{1},T_{1})\;>\;m_{2}(t_{1},T_{1})\;>0,$
where $A=(t_{1},T_{1})$ are the design variables that control the shape of the intermediate segment of the cure schedule. In practice, we explore a wide range of candidate points $A$ within the admissible bounds, construct the corresponding temperature profiles $T(\cdot;t_{1},T_{1})$, and pass each profile through DeepONet to evaluate $u(t_{3};T(\cdot;t_{1},T_{1}))$ and $DoC(t_{3};T(\cdot;t_{1},T_{1}))$. The constrained optimization problem is then solved over $A=(t_{1},T_{1})$ to identify a cure schedule that minimizes the final deformation while satisfying the degree-of-cure requirement.
Figure 13: Result of the cure schedule optimization. The left panel shows the optimized temperature profile with the selected intermediate point $A=(t_{1},T_{1})$, and the right panel shows the corresponding deformation history predicted by DeepONet for this temperature profile.
Figure 13 illustrates the performance of DeepONet for optimizing the deformation under the constrained optimization problem (4). The predicted optimal cure schedule corresponds to the intermediate point $A=(t_{1},T_{1})\approx(1.61\pm 0.02\penalty 10000\ \text{min},\,133.01\pm 0.03\penalty 10000\ ^{\circ}\text{C})$. It is worth noting that this optimized case is very close to the optimal cure profile previously obtained in [32].
8 Summary
We have presented a data-driven framework for predicting and mitigating process-induced deformation (PID) in composite laminates by combining high-fidelity simulations, targeted experiments, and deep operator learning. PID arises from mismatches between fiber and matrix responses during curing, driven by thermal expansion and cure shrinkage, and is modeled here using a two-mechanism thermo-chemical and stress–deformation framework calibrated and validated against manufacturing trials. Comparisons between measured and simulated PID show that the model captures the dominant physics, with baseline predictions lying within the experimental scatter and optimized cure cycles achieving an 8–10% reduction in deformation.
The validated model is then used to generate a diverse dataset of PID responses over a broad family of non-isothermal cure cycles. Using this dataset, we train a FiLM-enhanced DeepONet to predict the time histories of DoC, viscosity, and deformation for arbitrary temperature profiles, with Feature-wise Linear Modulation allowing the branch network to account for external parameters such as the initial DoC. To incorporate sparse experimental information, we apply transfer learning in which only the final layer of the pretrained network is updated using measured terminal deformation, while the underlying operator learned from simulation data is kept fixed.
Finally, an ensemble of DeepONet models is combined with Ensemble Kalman Inversion (EKI) to both quantify epistemic uncertainty and optimize the intermediate dwell point in the cure cycle, yielding temperature histories that minimize final deformation subject to a full-cure constraint. Overall, this framework illustrates how physics-based simulation, operator learning, and EKI-based uncertainty quantification can be integrated to construct reliable and data-efficient surrogates for composite cure process optimization, and it is readily extensible to other material systems as a foundation for robust cure-cycle design under uncertainty.
Acknowledgments
This work was supported as part of the AIM for Composites, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award #DE-SC0023389. This research used computing resources provided by the Center for Computation and Visualization (CCV) at Brown University.